# Degeneracy slopes, boundary slopes and exceptional surgery slopes

### v2

@@ Acknowledgments
The auhtor would like to thank Tsuyoshi Kobayashi and Yoa’v Rieck for helpful conversations. He also thanks to Steven Boyer for letting him know the reference [BGH24] regarding Theorem 2.
@@ 2. Distances between degeneracy slopes and boundary slopes
We first prepare basic terminology on essential laminations and essential branched surfaces. See [Cal03, Cal07] for details, and also [GO89, Definitions 1.1], [Gab97], and [Wu98].
A branched surface $B$ in a 3-manifold $M$ is defined as a 2-complex with a $C^{1}$ combing along the 1-skeleton, giving it a well-defined tangent space at every point, and generic singularities ([Cal07, Definition 6.16]).
@@
It is proved by [GO89, Proposition 4.5] that every essential lamination is fully carried by a branched surface with the following properties, which is called an essential branched surface.
@@
We remark that the surfaces considered in [Wu98] are assumed to be orientable, but the lemma holds for essential non-orientable surfaces as well. See the proof of [Wu98, Theorem 2.2].
@@
By [GO89, Proposition 4.5], there exists an essential branched surface $B$ that fully carries the lamination $\lambda$. After a small perturbation, we may assume that $B\cap F$ is a train track, say $\tau$.
@@
First, by Lemma 1, we obtain an essential branched surface, say $B_{1}$, from $B$ by splitting, and obtain an essential surface, say $F_{1}$, by isotopies such that $F_{1}\cap B_{1}$ is an essential train track $\tau_{1}$ on $F_{1}$, that is, $\tau_{1}$ has no complementary disks or monogons. We remark that in the procedure to obtain $\tau^{\prime}$ from $\tau$, the number of switches must decrease when $\tau$ is inessential. See [Wu98, Proof of Theorem 2.2 and Figure 2.1].
@@
Suppose otherwise; $S$ is inessential in $X_{1}$. Then, $S$ must be $\partial$-compressible in $X_{1}$. Since $S$ intersects with $\partial M$, and is not a vertical annulus, there exists a $\partial$-compressing disk for $S$ whose boundary consists of an arc $\alpha$ on $S$ and an arc on $\partial M$. Note that the arc $\alpha$ in $F_{1}$ separates connected components of $\tau_{1}$ on $F_{1}$, for $F_{1}$ is $\partial$-incompressible.
@@
We here claim that each rectangular region $R$ on $S$ bounded by two adjacent $\alpha_{i}$’s contains at least one cusp, i.e., an arc of $S\cap\partial_{v}N(B^{\prime})$. If $R$ contains $\alpha^{\prime}_{j}$’s, then, as observed above, it contains at least one cusp. Suppose otherwise that $R$ contains no $\alpha^{\prime}_{j}$’s, and suppose $R$ contains no cusps to induce a contradiction. Note that, by the minimality of $S\cap A$, there are no bigons on $\partial M$ bounded by arcs on $\partial S$ and $\partial A$. In this situation, by using $R$, we find another essential annulus $A^{\prime}$ in $X^{\prime}$ whose one boundary component lies on $\partial M$ and the other on $\partial_{h}N(B^{\prime})$, such that $A^{\prime}$ has a different boundary slope on $\partial M$ from that of $A$. However, this contradicts the uniqueness of the degeneracy slope for the very full essential lamination $\lambda$, which follows from the fact that $X^{\prime}$ is a one-holed ideal polygon bundle over $S^{1}$. See also [Gab97, Remark 8.6].
@@ 3.1. Degeneracy slopes for alternating knots
Let $K$ be a non-trivial prime alternating knot in $S^{3}$ with the exterior $E(K)$ in $S^{3}$. Then, by [Men84], $E(K)$ is a compact, orientable, irreducible, atoroidal 3-manifold with incompressible torus boundary. Thus, for a degeneracy slope $\delta$ of a very full essential lamination in $E(K)$, Theorem 1 is applicable to $E(K)$.
Suppose, for a contradiction, that $\delta$ is non-meridional. Then, since the meridional surgery on $K$ yields $S^{3}$, $\delta$ is integral, i.e., $\delta=p/1$ for some integer $p$, by [Wu98, Theorem 2.5].
Consider the checkerboard surfaces $F_{1}$, $F_{2}$ for the reduced alternating diagram of $K$. These are known to be essential by [Aum56, DR99]. Let $r_{1}$ and $r_{2}$ denote the boundary slopes of $F_{1}$ and $F_{2}$, which are both integral slopes. By simple calculations, we have
@@
Here, $\mathrm{cr}(K)$ denotes the minimal crossing number of $K$, which is actually equal to the number of crossings in the reduced alternating diagram. For example, see [IM09]. Hence, the following holds;
@@ 3.2. Bounds on boundary slopes
Since $K$ is hyperbolic, the exterior $E(K)$ is a compact, orientable, irreducible, atoroidal 3-manifold with incompressible torus boundary. Then, as stated in the Introduction, by [Mos96, Theorem A] and [Cal07, Theorem 6.47. (Gabai–Mosher)], the exterior $E(K)$ contains a very full essential lamination. We consider its degeneracy slope $\delta$ on the boundary $\partial E(K)$.
@@
Note that if $F$ is orientable, $|\partial F|\geq 2$ holds, since $\Sigma$ is an integral homology sphere. Thus, we obtain
@@ 3.3. Bounds on exceptional surgery slopes
As in the proof of Theorem 3, since $K$ is hyperbolic, the exterior $E(K)$ is a compact, orientable, irreducible, atoroidal 3-manifold with incompressible torus boundary. Thus, by [Mos96, Theorem A] and [Cal07, Theorem 6.47 (Gabai–Mosher)], the exterior $E(K)$ contains a very full essential lamination. We consider its degeneracy slope $\delta$ on the boundary $\partial E(K)$.
Let $\gamma=p/q$ be a non-trivial exceptional surgery slope for $K$. Then, by [Wu98, Theorem 2.5] together with the Geometrization of 3-manifolds, we have $\Delta(\gamma,\delta)\leq 2$.
@@
In particular, by considering the longitudinal slope, that is, $0/1$, of the minimal genus Seifert surface of genus $g(K)$, which is shown to be essential, we have

### v3

@@ Acknowledgments
The auhtor would like to thank Tsuyoshi Kobayashi and Yoa’v Rieck for helpful conversations. He also thanks to Steven Boyer for letting him know the reference [4] regarding Theorem 2 and the anonymous referee for a careful reading and helpful comments. This work was partially supported by JSPS KAKENHI Grant Number JP22K03301.
@@ 2. Distances between degeneracy slopes and boundary slopes
We first prepare basic terminology on essential laminations and essential branched surfaces. See [7, 8] for details, and also [13, Definitions 1.1], [14], and [48].
A branched surface $B$ in a 3-manifold $M$ is defined as a 2-complex with a $C^{1}$ combing along the 1-skeleton, giving it a well-defined tangent space at every point, and generic singularities ([8, Definition 6.16]).
@@
It is proved by [13, Proposition 4.5] that every essential lamination is fully carried by a branched surface with the following properties, which is called an essential branched surface.
@@
We remark that the surfaces considered in [48] are assumed to be orientable, but the lemma holds for essential non-orientable surfaces as well. See the proof of [48, Theorem 2.2].
@@
By [13, Proposition 4.5], there exists an essential branched surface $B$ that fully carries the lamination $\lambda$. After a small perturbation, we may assume that $B\cap F$ is a train track, say $\tau$.
@@
First, by Lemma 1, we obtain an essential branched surface, say $B_{1}$, from $B$ by splitting, and obtain an essential surface, say $F_{1}$, by isotopies such that $F_{1}\cap B_{1}$ is an essential train track $\tau_{1}$ on $F_{1}$, that is, $\tau_{1}$ has no complementary disks or monogons. We remark that in the procedure to obtain $\tau^{\prime}$ from $\tau$, the number of switches must decrease when $\tau$ is inessential. See [48, Proof of Theorem 2.2 and Figure 2.1].
@@
Suppose otherwise; $S$ is inessential in $X_{1}$. Then, $S$ must be $\partial$-compressible in $X_{1}$. Since $S$ intersects with $\partial M$, and is not a vertical annulus, there exists a $\partial$-compressing disk for $S$ whose boundary consists of an arc $\alpha$ on $S$ and an arc on $\partial M$. Note that, since $F_{1}$ is $\partial$-incompressible, the arc $\alpha$ cobounds a disk on $F_{1}$ together with an arc on $\partial F_{1}$, which may contain some connected components of $\tau_{1}$ on $F_{1}$.
@@
We here claim that each rectangular region $R$ on $S$ bounded by two adjacent $\alpha_{i}$’s contains at least one cusp, i.e., an arc of $S\cap\partial_{v}N(B^{\prime})$. If $R$ contains $\alpha^{\prime}_{j}$’s, then, as observed above, it contains at least one cusp. Suppose otherwise that $R$ contains no $\alpha^{\prime}_{j}$’s, and suppose $R$ contains no cusps to induce a contradiction. That is, $R\cap\partial N(B^{\prime})\subset\partial_{h}N(B^{\prime})$ holds. Note that, by the minimality of $S\cap A$, there are no bigons on $\partial M$ bounded by arcs on $\partial S$ and $\partial A$. In this situation, by using $R$, we find another essential annulus $A^{\prime}$ in $X^{\prime}$ whose one boundary component lies on $\partial M$ and the other on $\partial_{h}N(B^{\prime})$ as follows. Consider the arcs $R\cap A$ on $A$, which is a pair of essential arcs $\alpha_{i}$’s on $A$. They cobounds a rectangular region $R^{\prime}$ on $A$. (Actually there are two such regions, but either one can be used.) The union $R\cup R^{\prime}$ gives a vertical, thus essential, annulus $A^{\prime}$ in $X^{\prime}$ whose one boundary component lies on $\partial M$ and the other on $\partial_{h}N(B^{\prime})$, since $R\cap\partial N(B^{\prime})\subset\partial_{h}N(B^{\prime})$. By construction and the minimality of $S\cap A$, this $A^{\prime}$ has a different boundary slope on $\partial M$ from that of $A$. However, this contradicts the uniqueness of the degeneracy slope for the very full essential lamination $\lambda$, which follows from the fact that $X^{\prime}$ is a one-holed ideal polygon bundle over $S^{1}$. See also [14, Remark 8.6].
@@ 3.1. Degeneracy slopes for alternating knots
Let $K$ be a non-trivial prime alternating knot in $S^{3}$ with the exterior $E(K)$ in $S^{3}$. Then, by [35], $E(K)$ is a compact, orientable, irreducible, atoroidal 3-manifold with incompressible torus boundary. Thus, for a degeneracy slope $\delta$ of a very full essential lamination in $E(K)$, Theorem 1 is applicable to $E(K)$.
Suppose, for a contradiction, that $\delta$ is non-meridional. Then, since the meridional surgery on $K$ yields $S^{3}$, $\delta$ is integral, i.e., $\delta=p/1$ for some integer $p$, by [48, Theorem 1.6(1)] plus the first few lines from the proof of [48, Theorem 2.5].
Consider the checkerboard surfaces $F_{1}$, $F_{2}$ for the reduced alternating diagram of $K$. These are known to be essential by [1, 10]. Let $r_{1}$ and $r_{2}$ denote the boundary slopes of $F_{1}$ and $F_{2}$, which are both integral slopes. By simple calculations, we have
@@
Here, $\mathrm{cr}(K)$ denotes the minimal crossing number of $K$, which is actually equal to the number of crossings in the reduced alternating diagram. For example, see [23]. Hence, the following holds;
@@ 3.2. Bounds on boundary slopes
Since $K$ is hyperbolic, the exterior $E(K)$ is a compact, orientable, irreducible, atoroidal 3-manifold with incompressible torus boundary. Then, as stated in the Introduction, by [37, Theorem A] and [8, Theorem 6.47. (Gabai–Mosher)], the exterior $E(K)$ contains a very full essential lamination. We consider its degeneracy slope $\delta$ on the boundary $\partial E(K)$.
@@
Note that if $F$ is orientable, $|\partial F|\geq 2$ holds. Because the image of the boundary map $\partial:H_{2}(E(K),\partial E(K))\to H_{1}(\partial E(K))$ is a Lagrangian, that is, any two element in the image intersect algebraically zero times, hence when $F$ is orientable, we should have $\partial(F)=0\in H_{1}(\partial E(K))$ and so $|\partial F|$ is even.
Thus, we obtain
@@ 3.3. Bounds on exceptional surgery slopes
As in the proof of Theorem 3, since $K$ is hyperbolic, the exterior $E(K)$ is a compact, orientable, irreducible, atoroidal 3-manifold with incompressible torus boundary. Thus, by [37, Theorem A] and [8, Theorem 6.47 (Gabai–Mosher)], the exterior $E(K)$ contains a very full essential lamination. We consider its degeneracy slope $\delta$ on the boundary $\partial E(K)$.
Let $\gamma=p/q$ be a non-trivial exceptional surgery slope for $K$. Then, by [48, Theorem 2.5] together with the Geometrization of 3-manifolds, we have $\Delta(\gamma,\delta)\leq 2$.
@@
In particular, by considering the longitudinal slope, that is, $0/1$, of the minimal genus Seifert surface of genus $g(K)$, which is known to be essential, we have